OpenStudy (anonymous):

4(7x-1)=3(2x+1)-4(x+6)

OpenStudy (anonymous):

what is the question

OpenStudy (johnweldon1993):

I supposed we solve for 'x' 4(7x - 1) = 3(2x + 1) - 4(x + 6) let's distribute 28x - 4 = 6x + 3 - 4x - 24 now lets combine like terms 28x - 6x + 4x = ...? 3 + 4 - 24 =...?